(a) detevm/ve theimpulse response h(t) (b) use H3レ I H
Answer #1

Since from given circuit

Vo(t)= Vin(t),so h(t)= Vo(t)/Vin(t) h(t)= 1

Vo(t)= Vin(t)*h(t)

If Vin(t)= del(t)

Vo(t)= del(t)* h(t) = 1. Vo(t)= 1

if Vin(t)=8u(t)

Then Vo(t)= 8u(t)